# Animatable Gaussians: Learning Pose-dependent Gaussian Maps for High-fidelity Human Avatar Modeling

### v3

@@ 3.4 Pose Projection Strategy
Benefiting from the effective avatar representation, our method can reconstruct detailed human appearances under the training poses. However, given the inherently data-driven nature of learning-based avatars, addressing generalization to novel poses is also necessary and important. RAM-Avatar [12] trains a VAE to transform the testing pose into an in-distribution one. In this work, we propose to utilize Principal Component Analysis (PCA) to project a novel driving pose signal into the distribution of seen training poses for better generalization. Specifically, given a pose condition represented by posed position maps, we extract valid points and concatenate them as a vector $\mathbf{x}_{t}\in\mathbb{R}^{3M}$ ($M$ is the point number). The feature of each training frame composes a matrix $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{T}]$, where $T$ is the number of training frames. We perform PCA on $\mathbf{X}$, producing $N$ principal components $\mathbf{S}=[\mathbf{s}_{1},\cdots,\mathbf{s}_{N}]\in\mathbb{R}^{3M\times N}$ and standard deviation of each component $\sigma_{i}$. Given position maps derived by a novel driving pose, we project the corresponding feature $\mathbf{x}$ into the PCA space by
@@ 4 Experiments
Results. As shown in Fig. 1 and Fig. 5, our method can create realistic avatars with high-fidelity dynamic details from multi-view videos. More sequential results animated by challenging out-of-distribution poses [54] like sports and dancing can be found in the Supp. document and video.
@@
Dataset. We mainly utilize two public datasets for the experiments, including 3 sequences with 24 views from THuman4.0 dataset [106] and 5 sequences with 160 views from ActorsHQ dataset [28] (but we only use 47 full-body views for avatar modeling). For the comparison with AvatarReX [107], we also request their dataset including 2 sequences with 16 views. THuman4.0 and AvatarReX datasets also provide the SMPL-X [59] registrations. We fit SMPL-X for ActorsHQ dataset using [102]. We split each sequence as training and testing chunks, and the training chunk contains 1500 $\sim$ 3000 frames.
@@
Metric. We adopt Peak Signal-to-Noise Ratio (PSNR), Structure Similarity Index Measure (SSIM) [87], Learned Perceptual Image Patch Similarity (LPIPS) [101] and Frechet Inception Distance (FID) [24] for quantitative experiments.
4.1 Comparison
We compare our method with recent state-of-the-art works on avatar quality and generalization. These works include both body-only (TAVA [43], ARAH [86], SLRF [106], PoseVocab [45]) and full-body (AvatarReX [107]) avatars.
@@ 4.1 Comparison
Figure 6: Qualitative comparison with state-of-the-art body-only avatars on novel pose synthesis.
Figure 7: Qualitative comparison with AvatarReX on both training pose reconstruction (a) and novel view synthesis (b).
Body-only Avatars. We compare our method with TAVA, ARAH, SLRF, and PoseVocab on “subject00” and “subject02” sequences of THuman4.0 dataset [106]. We run the released codes of TAVA, ARAH and PoseVocab on the dataset, and request the results of SLRF from the authors. We present qualitative comparisons on novel pose synthesis in Fig. 6. In contrast to other methods, our approach excels in animating highly realistic avatars with significant improvement on high-fidelity dynamic details, including garment wrinkles, logos and other textural patterns. The quantitative comparison is also performed on the testing chunk of the “subject00” sequence as shown in Tab. 1, and these numerical results prove that our method achieves more accurate animation. Although PoseVocab and SLRF introduce a learnable pose dictionary or local NeRFs to improve the representation ability of the NeRF MLP, they still suffer from the low-frequency bias [80] of MLPs and fail to create highly realistic avatars. Contrarily, our method leverages powerful 2D CNNs and explicit 3D Gaussian splatting, thus achieving modeling finer-grained dynamic appearances.
@@
Table 1: Quantitative comparison with state-of-the-art body-only avatars.
@@
Table 2: Quantitative comparison with AvatarReX.
@@
Full-body Avatars. Full-body avatars including TotalSelfScan [13], X-Avatar [74] and AvatarReX [107] can realize expressive control of the body, hands and face. TotalSelfScan reconstructs full-body avatars from monocular self-rotation videos, and only displays animations that appear very rigid. X-Avatar requires 3D human scans under different poses as input for creating avatars. AvatarReX is the most relevant work with our method, i.e., creating avatars from multi-view videos. Fig. 7 shows the comparison with AvatarReX on both training and novel poses. Fig. 7 (a) demonstrates that our method can reconstruct more faithful and vivid details compared with AvatarReX. Although AvatarReX introduces local feature patches to encode more details, it remains constrained by the representation ability of the conditional NeRF MLPs. Fig. 7 (b) shows that given a novel pose, our method not only generates more realistic details but also produces more reasonable non-rigid deformation, particularly for long dresses, in comparison with AvatarReX. This is attributed to the ability of our method to learn pose-dependent deformations on a character-specific template that has already modeled the basic shape of the wearing garments. In contrast, AvatarReX learns sparse node translations on the naked SMPL model, resulting in artifacts for long dresses. Tab. 2 reports the quantitative comparison on training pose reconstruction. Our method also outperforms AvatarReX on the reconstruction accuracy.
Animation Speed. We additionally compare our method with other works on the animation speed in Tab. 3. Benefiting from the efficient 3D Gaussian splatting [35], our method can realize fast inference for animation.
Conclusion. Overall, benefiting from the powerful 2D CNNs and explicit 3D Gaussian splatting, our method, Animatable Gaussians, achieves lifelike avatar modeling with more dynamic, realistic and generalized appearances in comparison to state-of-the-art approaches.
Table 3: Comparison on animation speed. These framerates are evaluated on a PC with one RTX 3090 when rendering images at a resolution of $1024\times 1024$. We highlight the highest and second-highest framerates.
Method TAVA [43] ARAH [86] SLRF [106] PoseVocab [45]
AvatarReX [107]
AvatarReX
Ours
Framerate (FPS) $\uparrow$ 0.003 0.07 0.16 0.20 0.03 25 10
@@
4.2 Ablation Study
We qualitatively evaluate the core contributions of our method in this subsection. Quantitative results and additional experiments can be found in the Supp. document.
@@ 4.2 Ablation Study
Parametric Template. We evaluate the learned parametric template by replacing it with a naked parametric model, SMPL-X [59]. Fig. 8 shows that SMPL-X fails to represent the long dress whose topology is not consistent with the SMPL-X model, yielding poor generalization to novel poses. Conversely, our character-specific template is adaptively reconstructed from the input video to model the basic shape of the wearing garments.
Backbones. To demonstrate the superior representation ability of 2D CNNs (StyleUNet in our settings), we replace StyleUNet with a coordinate-based MLP and a standard U-Net [70], respectively. The MLP takes a canonical point and pose vector as input, and returns the 3D Gaussian attributes of this point. While the standard U-Net replaces the StyleUNet as the backbone. Fig. 9 shows the animation results of our method with StyleUNet and the baselines with MLPs and U-Net, respectively. First, it demonstrates that 2D CNNs are able to regress more detailed and realistic appearances, while MLPs suffer from limited representation ability, yielding blurry animation results. Second, StyleUNet outperforms the standard U-Net because of the additional modules (including style modulation and “To/From-RGB” modules) inherited from StyleGAN [32]. Overall, the 2D parameterization and StyleUNet enable our method to model high-quality human dynamic appearances.
Pose Projection. We evaluate the pose projection strategy by removing it, i.e., directly inputting the position map into the StyleUNet. Fig. 10 shows the animation results with and without the pose projection under novel poses, respectively. It demonstrates that direct extrapolation with the novel position map results in unreasonable 3D Gaussians, since no similar poses in the training dataset. In contrast, the pose projection guarantees that the reconstructed position maps (Eq. 8) lie within the distribution of training poses, leading to reasonable and vivid synthesized appearances.
5 Discussion
Conclusion. We present Animatable Gaussians, a new avatar representation for creating lifelike human avatars with highly dynamic, realistic and generalized appearances from multi-view RGB videos. Compared with implicit NeRF-based approaches, we introduce the explicit point-based representation, 3D Gaussian splatting, into the avatar modeling, and leverage powerful 2D CNNs for modeling higher-fidelity human appearances. Based on the proposed template-guided parameterization and pose projection strategy, our method can not only faithfully reconstruct detailed human appearances, but also generate realistic garment dynamics for novel pose synthesis. Overall, our method outperforms other state-of-the-art avatar approaches, and we believe that the proposed 3D Gaussian splatting-based avatar representation will make progress towards effective and efficient 3D human representations.
Limitation. Our method entangles the modeling of the human body and clothes, limiting to changing the clothes of the avatar for applications like virtual try-on. A possible solution is to separately represent the body and clothes with multi-layer 3D Gaussians as NeRF-based approaches [15, 16]. Moreover, our method relies on the multi-view input to reconstruct a parametric template, limiting the application for modeling loose clothes from a monocular video.
Potential Social Impact. Our method can synthesize virtual motions of the human avatar to generate fake videos, so we must carefully employ this technology.
Acknowledgement. This paper is supported by National Key R&D Program of China (2022YFF0902200), the NSFC project No.62125107 and the Beijing Municipal Science & Technology Z231100005923030.
Supplemental Document
Figure A: Example sequential animation results by our method. Each row is an animation sequence involving 3 subjects. Our method can generate realistic and reasonable dynamic details even under novel poses from the AMASS dataset [54].
In this supplemental document, we will show implementation & experiment details, more results and additional experiments.
Appendix A Implementation Details
Template Reconstruction. We optimize an SDF and color field represented by an MLP consisting of intermediate layers with (512, 256, 256, 256, 256, 256) neurons. Given a posed point, we find accurate correspondence in the canonical space by root finding. Following ARAH [86], we initialize the correspondence as the canonical position that is computed by inverse skinning based on blending weights of the closest SMPL vertex. Different from SNARF [7] and ARAH [86] that utilize the Broyden’s method [5] to solve Eq. 3, we employ the Gauss-Newton method by implementing a customized CUDA kernel. The training loss of template reconstruction involves an RGB loss, a mask loss and an Eikonal loss [17].
Network Architecture. The network in our avatar representation is composed of StyleUNet [83], a conditional StyleGAN-based [32] generator. Differently, we adapt the original StyleUNet by incorporating two decoders to predict both front & back Gaussian maps. The resolution of the input position map is $512\times 512$, and the resolution of the output Gaussian maps is $1024\times 1024$. Specifically, we utilize three different StyleUNets to output color (3-channel), position (3-channel) and other Gaussian attributes (8-channel). In the color StyleUNet, we modulate the color output with a view direction map to model view-dependent effects. Each pixel on the view direction map indicates the angle between the view direction and the template normal. The view direction map is encoded through a tiny CNN, then the encoded feature map is injected into an intermediate decoder layer of the color StyleUNet.
Training. We adopt the Adam optimizer [37] for training the StyleUNet with a learning rate of $5\times 10^{-4}$. The loss weights are set as: $\lambda_{\text{perceptual}}=0.1$, $\lambda_{\text{reg}}=0.005$. The batch size is 1, the total iteration number is 500k, and the training procedure takes about two days on one RTX 4090.
Appendix B Experiment Details
Metric Evaluation. We utilize PSNR, SSIM [87], LPIPS [101] and FID [24] as the metrics for quantitative evaluations. PSNR and SSIM are computed on the entire image at the original resolution, while LPIPS and FID are computed on the cropped minimal square that covers the human body.
Comparison with Body-only Avatars. The quantitative comparison (Tab. 1 in the main paper) is conducted on the “subject00” sequence of THuman4.0 dataset [106]. The first 2000 frames are utilized as the training dataset, and the numerical results of Tab. 1 are evaluated on the rest 500 frames and the “cam18” camera view.
Comparison with Full-body Avatars. The quantitative comparison (Tab. 2 in the main paper) is conducted on the “avatarrex_zzr” sequence of AvatarReX dataset [107]. The numerical results of Tab. 2 are evaluated on the first 500 frames and the “22010710” camera view.
Appendix C Additional Sequential Results
Fig. A shows additional sequential results animated by challenging out-of-distribution pose sequences from the AMASS dataset [54], including basketball, football and dancing. It demonstrates that our method can also generate realistic and reasonable dynamic details under out-of-distribution poses thanks to the effective avatar representation and pose projection strategy.
Appendix D Additional Experiments
Table A: Quantitative ablation study on the parametric template.
@@ Appendix D Additional Experiments
Table B: Quantitative comparison between representations with different backbones.
@@
Table C: Quantitative ablation study on pose projection.

### v4

@@ III-E1 Pose Projection Strategy
Benefiting from the effective avatar representation, our method can reconstruct detailed human appearances under the training poses. However, given the inherently data-driven nature of learning-based avatars, addressing generalization to novel poses is also necessary and important. RAM-Avatar [115] trains a VAE to transform the testing pose into an in-distribution one. In this work, we propose to utilize Principal Component Analysis (PCA) to project a novel driving pose signal into the distribution of seen training poses for better generalization. Specifically, given a pose condition represented by posed position maps, we extract valid points and concatenate them as a vector $\mathbf{x}_{t}\in\mathbb{R}^{3M}$ ($M$ is the point number). The vector of each training frame composes a matrix $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{T}]$, where $T$ is the number of training frames. We perform PCA on $\mathbf{X}$, producing $N$ principal components $\mathbf{S}=[\mathbf{s}_{1},\cdots,\mathbf{s}_{N}]\in\mathbb{R}^{3M\times N}$ and standard deviation of each component $\sigma_{i}$. Given position maps derived by a novel driving pose, we project the corresponding vector $\mathbf{x}$ into the PCA space by
@@
III-E2 Relighting
Thanks to the introduction of PBR into the avatar representation, our method is able to relight the avatar under novel illumination. Specifically, given a novel pose, we first feed it into the avatar networks to predict Gaussian, normal, albedo and roughness maps. Note that we do not predict light visibility maps during testing because we empirically find it cannot accurately generalize to novel poses. We hypothesize the reasons are that the visibility is heavily determined by global self-occlusions, and the variety of training poses is limited. To this end, we directly calculate the visibility of posed 3D Gaussians via point-based ray tracing [20]. Then, given a novel environment map, we inject it with the predicted normal, abledo, and roughness as well as calculated visibility into the PBR equation (Eq. 10) to compute the PBR color of each 3D Gaussian. Finally, the relit image is obtained by splatting-based rasterization. As illustrated in Fig. 7, our method can produce photorealistic animation under novel illumination.
@@ IV-A1 Animation
As shown in Fig. 1 and Fig. 5, our method can create realistic avatars with high-fidelity dynamic details from multi-view videos. We also show results animated by challenging out-of-distribution poses from AMASS dataset [116] in Fig. 6. Thanks to the effective avatar representation and pose projection strategy, our method can produce animation results with highly dynamic, realistic and generalized appearances.
Figure 6: Example sequential animation results by our method. Each row is an animation sequence involving 3 subjects. Our method can generate realistic and reasonable dynamic details even under novel poses from the AMASS dataset [116].
Figure 7: Example intrinsic decomposition and relighting results under novel poses by our method. The first four columns display the decomposition results of material and geometry. The fifth column showcases the relighting outcomes with a directional point light, whereas the remaining columns are relit using real environment maps.
IV-A2 Relighting
In addition to animation, our method extends support to relighting applications. By naturally incorporating physically-based inverse rendering techniques into our powerful avatar representation, we achieve accurate intrinsic decomposition and vivid relighting results, as demonstrated in Fig. 1 and Fig. 7.
Please refer to the supplementary video for more sequential animation and relighting results.
IV-B Dataset and Metric
@@
We mainly utilize three public datasets for the experiments, including 3 sequences with 24 views from THuman4.0 dataset [7], 3 sequences with 16 views from AvatarReX dataset [66] and 5 sequences with 160 views from ActorsHQ dataset [117] (we only use 47 full-body views for avatar modeling). THuman4.0 and AvatarReX datasets also provide the SMPL-X [110] registrations. We fit SMPL-X for ActorsHQ dataset using the method proposed by Zhang et al.[118]. We split each sequence as training and testing chunks, and the training chunk contains 1500 $\sim$ 3000 frames.
@@
We adopt Peak Signal-to-Noise Ratio (PSNR), Structure Similarity Index Measure (SSIM) [119], Learned Perceptual Image Patch Similarity (LPIPS) [113] and Fréchet Inception Distance (FID) [120] for quantitative experiments. PSNR and SSIM are computed on the entire image at the original resolution, while LPIPS and FID are computed on the cropped minimal square that covers the human body.
@@
Figure 8: Qualitative comparison with state-of-the-art NeRF-based body-only avatars including TAVA [8], ARAH [59], SLRF [7] and PoseVocab [63] on novel pose synthesis.
TABLE I: Loss weights in the training process.
Loss Weight Value
$\lambda_{\text{perceptual}}$ 0.1
$\lambda_{\text{normal}}$ 0.2
$\lambda_{\text{vis}}$ 0.1
$\lambda_{\text{geo}}$ 0.01
$\lambda_{\text{albedo}}$ 0.005
$\lambda_{\text{roughness}}$ 0.005
IV-C Implementation Details
IV-C1 Template Reconstruction
We optimize an SDF and color field represented by an MLP consisting of intermediate layers with (512, 256, 256, 256, 256, 256) neurons. Given a posed point, we find accurate correspondence in the canonical space by root finding. Following ARAH [59], we initialize the correspondence as the canonical position that is computed by inverse skinning based on blending weights of the closest SMPL vertex. Different from SNARF [44] and ARAH [59] that utilize the Broyden’s method [121] to solve Eq. 3, we employ the Gauss-Newton method by implementing a customized CUDA kernel. The training loss of template reconstruction involves an RGB loss, a mask loss and an Eikonal loss [122].
IV-C2 Network Architecture
The network in our avatar representation is composed of StyleUNet [17], a conditional StyleGAN-based [14] generator. Differently, we adapt the original StyleUNet by incorporating two decoders to predict both front & back Gaussian and intrinsic maps. The resolution of the input position map is $512\times 512$, and the resolution of the output Gaussian and intrinsic maps is $1024\times 1024$. Specifically, we utilize five different StyleUNets to output color (3-channel), position (3-channel), other Gaussian attributes (8-channel), albedo & roughness (4-channel) and normal & visibility (19-channel). The visibility is represented as 16-dimensional SH coefficients. In the color StyleUNet, we modulate the color output with a view direction map to model view-dependent effects. Each pixel on the view direction map indicates the angle between the view direction and the template normal. The view direction map is encoded through a tiny CNN, then the encoded feature map is injected into an intermediate decoder layer of the color StyleUNet.
IV-C3 Training
We adopt the Adam optimizer [123] for training the network with a learning rate of $5\times 10^{-4}$. The loss weights are set as illustrated in Tab. I. The batch size is 1, the total iteration number is 500k, and the training procedure takes about two days on one RTX 4090.
IV-C4 Running Time
For only animation, it takes around 0.13 secs to render one frame. Given novel illumination for relighting, it takes $4\sim 10$ secs to synthesize one frame due to the additional computational cost of light visibility computation in PBR. The relighting time cost mainly depends on the numbers of 3D Gaussians and sampled rays.
@@ IV-C4 Running Time
TABLE II: Quantitative comparison with state-of-the-art NeRF-based body-only avatars.
@@
IV-D Comparison on Avatar Animation
In the comparisons with state-of-the-art animatable avatars, we first compare our method with NeRF-based approaches including both body-only (TAVA [8], ARAH [59], SLRF [7], PoseVocab [63]) and full-body (AvatarReX [66]) avatars. Then we compare our method with concurrent 3D Gaussian splatting-based avatars including 3DGS-Avatar [91] and GaussianAvatar [95].
IV-D1 NeRF-based Body-only Avatars
We compare our method with TAVA, ARAH, SLRF, and PoseVocab on “subject00” and “subject02” sequences of THuman4.0 dataset [7]. We run the released codes of TAVA, ARAH and PoseVocab on the dataset, and request the results of SLRF from the authors. We present qualitative comparisons on novel pose synthesis in Fig. 8. In contrast to other methods, our approach excels in animating highly realistic avatars with significant improvement on high-fidelity dynamic details, including garment wrinkles, logos and other textural patterns. The quantitative comparison is also performed on the testing chunk (the 2000-2500 frames and “cam18” view) of the “subject00” sequence as shown in Tab. II, and these numerical results prove that our method achieves more accurate animation. Although PoseVocab and SLRF introduce a learnable pose dictionary or local NeRFs to improve the representation ability of the NeRF MLP, they still suffer from the low-frequency bias [9] of MLPs and fail to create highly realistic avatars. Contrarily, our method leverages powerful 2D CNNs and explicit 3D Gaussian splatting, thus achieving modeling finer-grained dynamic appearances.
Figure 9: Qualitative comparison with AvatarReX [66] on both training pose reconstruction (a) and novel view synthesis (b).
TABLE III: Quantitative comparison with the state-of-the-art NeRF-based full-body avatar.
@@ IV-D2 NeRF-based Full-body Avatars
Full-body avatars including TotalSelfScan [64], X-Avatar [65] and AvatarReX [66] can realize expressive control of the body, hands and face. TotalSelfScan reconstructs full-body avatars from monocular self-rotation videos, and only displays animations that appear very rigid. X-Avatar requires 3D human scans under different poses as input for creating avatars. AvatarReX is the most relevant work with our method, i.e., creating avatars from multi-view videos. Fig. 9 shows the comparison with AvatarReX on both training and novel poses. Fig. 9 (a) demonstrates that our method can reconstruct more faithful and vivid details compared with AvatarReX. Although AvatarReX introduces local feature patches to encode more details, it remains constrained by the representation ability of the conditional NeRF MLPs. Fig. 9 (b) shows that given a novel pose, our method not only generates more realistic details but also produces more reasonable non-rigid deformation, particularly for long dresses, in comparison with AvatarReX. This is attributed to the ability of our method to learn pose-dependent deformations on a character-specific template that has already modeled the basic shape of the wearing garments. In contrast, AvatarReX learns sparse node translations on the naked SMPL model, resulting in artifacts for long dresses. Tab. III reports the quantitative comparison on training pose reconstruction. Our method also outperforms AvatarReX on the reconstruction accuracy. The numerical results are evaluated on the first 500 frames and the “22010710” camera view in the “avatarrex_zzr” sequence from AvatarReX dataset.
Figure 10: Qualitative comparison with state-of-the-art 3D Gaussian splatting-based avatars including 3DGS-Avatar [91] and GaussianAvatar [95].
TABLE IV: Quantitative comparison with state-of-the-art 3D Gaussian splatting-based avatars.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
Ours 30.3607 0.9682 0.0339 33.4665
3DGS-Avatar [91] 28.7836 0.9511 0.0418 49.3673
GaussianAvatar [95] 26.9497 0.9389 0.0407 38.5387
Figure 11: Qualitative comparison with the state-of-the-art human performance relighting methods, R4D [104] and RA [107].
IV-D3 3D Gaussian Splatting-based Avatars
We qualitatively compare our method with 3DGS-Avatar [91] and GaussianAvatar [95] using “Actor01” and “Actor02” sequences from ActorsHQ dataset [117] in Fig. 10. It shows that our method outperforms other approaches by a large margin on the avatar quality, especially the dynamic wrinkles of the garments. 3DGS-Avatar [91] utilizes pure MLPs to regress the non-rigid deformations and pose-dependent appearances, suffering from the limited capacity of MLPs. Although GaussianAvatar [95] employs a 2D U-Net [124] to regress Gaussian parameters on SMPL UV space, it freezes the opacity and rotation attribute as pose-agnostic variables, deteriorating the modeling ability of the whole model. Moreover, these two methods utilize the naked SMPL model to parameterize 3D Gaussians and both fail to model detailed motions and appearances of loose clothes as shown in the top row of Fig. 10. We also report the numerical results in Tab. IV, and our method also quantitatively outperforms 3DGS-Avatar and GaussianAvatar. The numerical results are computed on the 48-548 frames and the “Cam127” camera view in the “Actor01/Sequence1” from ActorsHQ dataset.
TABLE V: Quantitative comparison with state-of-the-art human performance relighting method.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
Ours 31.6339 0.9836 0.0208 24.1962
RA[107] 24.8013 0.9694 0.0738 57.3957
R4D[104] 23.0883 0.9608 0.0968 112.9600
IV-E Comparison on Human Performance Relighting
We compare our method with the state-of-the-art human performance relighting methods Relighting4D (R4D) [104] and Relightable and Animatable Avatar (RA) [107] on ActorsHQ [117] and AvatarReX [66] datasets.
We present the qualitative and quantitative comparison in Fig. 11 and Tab. V. Our method excels in precise intrinsic decomposition due to the proposed physically-based avatar representation, particularly evident in the albedo and normal details. This highlights the enhanced capabilities of our approach in capturing intricate details in both geometry and appearance. Moreover, the accurate material and geometry can then contribute to natural relighting under novel illumination. However, R4D [104] and RA [107], as NeRF-based implicit methods, fail to model detailed material and geometry of the character due the limited capacity of their avatar representations. These two approaches produce blurry results on both texture and geometry, leading to diminished relighting results.
IV-F Ablation Study
We evaluate the core contributions of our method in this subsection.
@@ IV-F Ablation Study
IV-F1 Parametric Template
@@
We evaluate the learned parametric template by replacing it with a naked parametric model, SMPL-X [110]. Fig. 12 shows that SMPL-X fails to represent the long dress whose topology is not consistent with the SMPL-X model, yielding poor generalization to novel poses. Conversely, our character-specific template is adaptively reconstructed from the input video to model the basic shape of the wearing garments. We also quantitatively compare the reconstructed parametric template with the naked SMPL-X model [110] on the animation accuracy in Tab. VI. It shows that the reconstructed template can animate the 3D Gaussians more accurately.
TABLE VI: Quantitative ablation study on the parametric template.
@@
TABLE VII: Quantitative comparison between representations with different backbones.
@@
IV-F2 Backbones
To demonstrate the superior representation ability of 2D CNNs (StyleUNet in our settings), we replace StyleUNet with a coordinate-based MLP and a standard U-Net [124], respectively. The MLP takes a canonical point and pose vector as input, and returns the 3D Gaussian attributes of this point. While the standard U-Net replaces the StyleUNet as the backbone. Fig. 13 and Tab. VII show the qualitative and quantitative animation results of our method with StyleUNet and the baselines with MLPs and U-Net, respectively. First, it demonstrates that 2D CNNs are able to regress more detailed and realistic appearances, while MLPs suffer from limited representation ability, yielding blurry animation results. Second, StyleUNet outperforms the standard U-Net because of the additional modules (including style modulation and “To/From-RGB” modules) inherited from StyleGAN [14]. Overall, the 2D parameterization and StyleUNet enable our method to model high-quality human dynamic appearances.
TABLE VIII: Quantitative ablation study on pose projection.